# http://www.michaeltimothybennett.com/

### v2

@@ Abstract
Simplicity is held by many to be the key to general intelligence. Simpler models tend to “generalise”, identifying the cause or generator of data with greater sample efficiency. The implications of the correlation between simplicity and generalisation extend far beyond computer science, addressing questions of physics and even biology. Yet simplicity is a property of form, while generalisation is of function. In interactive settings, any correlation between the two depends on interpretation. In theory there could be no correlation and yet in practice, there is. Previous theoretical work showed generalisation to be a consequence of “weak” constraints implied by function, not form. Experiments demonstrated choosing weak constraints over simple forms yielded a $110-500\%$ improvement in generalisation rate. Here we measure the complexity of weak constraints, and show that if one does not presuppose an abstraction layer, then all have equal complexity. However, in the context of a spatially and temporally extended abstraction layer, efficiency demands weak constraints take simple forms, and simplicity becomes correlated with generalisation. Simplicity has no causal influence on generalisation, but appears to due to confounding.
Keywords:
complexity weakness causality AGI information theory.
@@ 1 Introduction
Complexity is a quality of systems we find difficult to understand. Formal analogues include entropy [1], compression [2, 3] and even fractal dimension [4]. Physicist Leonard Susskind believes complexity may be the key to a unified theory of physics [5, 6]. Cyberneticist Francis Heylighen recently argued that goals are attractors of dynamical systems that self organise in complex reaction networks [7]. If complex reaction networks do self organise as he argues [8], then it goes some way towards explaining the origins of goal directed behaviour, and thus life. Finally, many hold that complexity is the key to general intelligence. Language models like GPT-4 amount to compressed representations of human language [9]. Simpler objects can be compressed to greater extents, because they exhibit self similarity. As a result, some hold that compression is general intelligence, meaning a general reinforcement learning agent like AIXI [10] can use Solomonoff Induction [11] to maximise expected reward across a wide range of environments. Yet for all that complexity seems to be at the heart of every matter, it has profound flaws. As a qualitative indicator of how subjectively difficult a system might be to understand, it makes perfect sense. It makes far less as an indicator of anything objective. Ockham’s Razor, the epistemic principle that simpler statements are more likely to hold true, can be understood as the claim that our subjective perceptions of complexity reflect an objective property of our environment. There is no obvious reason this should be the case, and yet it is [12]. Simpler statements tend to be more accurate representations of reality. The aforementioned Solomonoff Induction formalises Ockham’s Razor, meaning AIXI is based on the premise that simpler models are more accurate representations of the environment. AIXI is a superintelligence in the sense that it maximises Legg-Hutter intelligence [13]. However, Jan Leike later showed that any claim regarding AIXI’s performance is “entirely subjective”. Legg-Hutter intelligence is measured with respect to a fixed Universal Turing Machine (UTM), and AIXI is only optimal if it uses exactly the same UTM [14]. This calls into question the viability of complexity based induction systems in interactive settings. Their performance is subjective, and from a pragmatic standpoint it is only objective performance claims that matter. Leike’s result suggests there could be no correlation between objective performance and subjective complexity. This concurs with what seems intuitively obvious, that complexity is an aspect of interpretation. Complexity is a measure of form, not function. So why does the subjective perception of simplicity tend to correlate with objective performance?
@@ 1.0.1 What exactly is complexity supposed to indicate?:
As it is used in AGI research [10], complexity is intended to help us infer the program which generated or caused data. If one can identify that which caused past data, then one can “generalise” to predict the outcomes in future interactions, to maximise performance [15]. We are concerned with adaptation or “the ability to generalise” [16], not any specific circumstance. This is because any system, even a rote learner, can eventually identify cause given enough data and memory (by simply rote learning every outcome until it has a complete behavioural specification of the causal program). So assuming one can correctly infer cause, then the amount of data one requires to do so is the sole measure of performance. We refer to this as sample efficiency. The more sample efficiently one can infer cause, the greater one’s ability to generalise and adapt to any desired end. Thus, we take intelligence to be a measure of the sample efficiency in generalisation.
@@ 1.0.2 Key questions:
We build upon previous work [17, 15], in which:
@@
    Maximising simplicity of policies was proven unnecessary and insufficient to maximise sample efficiency [17, prop. 3].
@@
    Maximising policy “weakness” was proven necessary and sufficient to maximise sample efficiency [17, prop. 1,2] and identify cause [15]. In experiments, weak policies outperformed simple by $110-500\%$.
@@ 2 The Formalism
The following definitions are shared with [18, 19], which apply them to biological and philosophical perspectives.
@@
  • We assume an infinite set $\Phi$ whose elements we call states.
@@
  • We say a statement is true iff it is an aspect of the environment realised by the present state.
@@ 2.0.1 Learning and inference definitions:
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to alpha is less than the probability that $l_{2}$ generalizes to $\alpha$.
@@
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [17]).
@@ 3 Arguments and Results
The distinction between software and hardware, though useful, is unsuitable for reasoning about cause [15]. A software agent in an interactive setting performs subjectively, relying on hardware through which it is interpreted [14]. If we are to understand complexity, we must understand what the concept entails in the absence of such “abstraction layers”. We must ascertain what is objective rather than subjective, and so we must begin at the level of the environment rather than the agent. To this end, previous work [17, 15, 20] proposed a “de facto” pancomputational [21] model of all possible environments and aspects thereof, including all abstractions and goal directed behaviours, which we refine and extend. While the formalism we present here is a departure from past work, it is equivalent with respect to those previously published proofs which we reference [17]. The formalism is not computational in the sense of relying on symbols and quantities, or any other high level abstraction interpreted by a human mind. The assumptions we make are extremely weak.
Axiom 1: There are things which exist, and things which do not. We call the things which exist the environment.
Axiom 2: The environment has at least one state. To say it has more than one state is to claim there exists at least one “dimension”. A dimension is a set of points, for example time, we do not make the additional claim that there is any order to that set. Each state is the environment at a different point. Each state is “extended” along dimensions, meaning no two states can occupy the same points in all dimensions (in other words, states are like the environment perceived from different positions in time or some other dimensions by an omniscient observer).
These are all the assumptions we need for our version of pancomputationalism. We don’t even need to speculate about any internal structure states might have. States differ along one or more dimensions. Any “fact” about a state’s internal structure can be defined by its relation to other states. The existence of sets is implied by the existence of states (because there is more than one “thing” that exists), and a fact is just the set of states in which it holds (the truth conditions of a “declarative program”).
@@
As truth is defined in existential terms, there is nothing which is not a fact of the environment, and no environment which we cannot describe as a set of facts. In other words, this is a minimalist formalism of everything. These are not statements evaluated by a mind, and truth is not semantic in that sense. An environment could be like our own, or not. It could be deterministic, in which case states follow a sequence, or non-deterministic, in which case they don’t [^4] In any case, the difference between deterministic and non-deterministic seems meaningless when you consider that DFAs and NFAs are equivalent [22].. It could be fantastical with magic and true names. It could even be a world constructed through the subjective experience of its inhabitants. All that matters is that an environment has states differentiated along one or more dimension, and from there we obtain the set of all facts. This gives us some notion of internal structure. An aspect of a state is just a set of facts about that state. With aspects in hand, we can define abstraction. An aspect is akin to a logical statement. It has a truth value given a state. We also define its extension (def. 2), which is all other aspects of which it is a part. This implies a heirarchy or “lattice” of aspects. Intuitively, an abstraction layer is like a window through which one can view part of the environment. A laptop computer could function as an abstraction layer, as could all or part of the system in which an embodied and embedded organism enacts cognition [23]. In precise terms an abstraction layer is implied by a vocabulary, which is a set of declarative programs. A vocabulary implies a formal language whose rules are determined by relations between states. An abstraction layer implies a set of “$\mathfrak{v}$-tasks” (def. 3), each of which is behaviour that defines a system. The formalisation of policies as causal intermediaries between inputs and outputs [24] then develops this into a causal depiction of goal directed behaviour (assuming a policy is implied by the inputs and outputs). For example, an organism could be a policy for a $\mathfrak{v}$-task which is behaviour that organism might enact. Again, we must emphasise this is a first principles approach. We do not assume symbols and Turing machines. Inputs, outputs and policies are all just sets of declarative programs. Whether something is goal directed is determined by the relations between states. The environment makes only one sort of value judgement (existence or non-existence), and is otherwise impartial. Goal directed behaviour is a value judgement, so we formalise this impartiality as a uniform distribution over tasks for the purposes of learning. Finally, we must now define complexity. The claims we make in the rest of this paper pertain to this notion of complexity.
@@
Complexity of Extension: If two statements have the same extension, then the simpler is the smaller set (def. 5). The complexity of an extension is the simplest statement of which it is the extension. This is analogous to other measures of complexity [2, 3], but facilitates comparison of abstraction layers.
@@ 3.1 Implications for Complexity
In accord with definition 2, the absence of abstraction means the vocabulary is the set of all declarative programs, meaning $\mathfrak{v}=P$. It follows that for every $l\in L_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ such the $\bigcap l=f$. Statements $l$ and $\{f\}$ are equivalent iff $E_{l}=E_{\{f\}}$, which is exactly the case here because $\bigcap l=f$. [17, prop. 1,2] shows maximising weakness is necessary and sufficient to maximise the probability of generalisation, which means weakness maximises sample efficiency (is the optimal proxy). This means sample efficiency is determined by the cardinality of extension. For every correct policy $l$ of every task in $\Gamma_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$. Policy complexity can be minimised regardless weakness, because the simplest representation of every extension is simplicity of length $1$.
In this sense, complexity is an illusion created by abstraction. In the absence of any particular abstraction, all behaviours (extensions) are implied by statements of the same complexity. However, that doesn’t change the fact that simpler forms have been observed to generalise better. This raises the question; what additional assumptions can we make that would explain the correlation?
@@ 3.1.1 Time, space and causal confounding:
We now make the additional assumption that vocabularies are finite. Every aspect of the world in which we exist appears to be spatially and temporally extended, so for the sake of understanding complexity we assume this is true of all environments [^5] Something is spatially and temporally extended if it occupies a location in space and time, and nothing else can occupy that same location.. We hold that this justifies the assumption of a finite vocabulary, because in our spatially and temporally extended environment the amount of information which can influence a system in a finite area of time and space is finite [25].
@@
We already have that policy weakness causes sample efficiency, in that it is necessary and sufficient to maximise it in order to maximise sample efficiency. Continuing from proof 1, in a finite vocabulary, there may not exist $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$, which means the complexity of all extensions will not be the same. If we choose any vocabulary in which weaker aspects take simpler forms, then simplicity will be correlated with weakness and so will also be correlated with sample efficiency. This means we would choose $\mathfrak{v}$ s.t. for all $a,b\in L_{\mathfrak{v}}$, the simpler statement has the larger extension, meaning $a<_{w}b\leftrightarrow a<_{d}b$. For example, suppose $P=\{a,b,c…\},\ a=\{1,2,4\},\ b=\{1,3,4\},\ \mathfrak{v}=\{a,b\},\ L_{\mathfrak{v}}=\{\{a\},\{b\},\{a,b\}\}$, then it follows $\{a,b\}<_{w}\{a\},\ \{a,b\}<_{w}\{b\},\ \{a,b\}<_{d}\{a\},\ \{a,b\}<_{d}\{b\}$.

### v3

@@ Abstract
Simplicity is held by many to be the key to general intelligence. Simpler models tend to “generalise”, identifying the cause or generator of data with greater sample efficiency. The implications of the correlation between simplicity and generalisation extend far beyond computer science, addressing questions of physics and even biology. Yet simplicity is a property of form, while generalisation is of function. In interactive settings, any correlation between the two depends on interpretation. In theory there could be no correlation and yet in practice, there is. Previous theoretical work showed generalisation to be a consequence of “weak” constraints on implied by function, not form. Experiments demonstrated choosing weak constraints over simple forms yielded a $110-500\%$ improvement in generalisation rate. Here we show that all constraints can take equally simple forms, regardless of weakness. However if forms are spatially extended, then function is represented using a finite subset of forms. If function is represented using a finite subset of forms, then we can force a correlation between simplicity and generalisation by making weak constraints take simple forms. If function determined by a goal directed process (e.g. natural selection), then efficiency demands weak constraints take simple forms. Complexity has no causal influence on generalisation, but appears to due to confounding.
@@ 1 Introduction
Complexity is a quality of systems we find difficult to understand. Formal analogues include entropy [sep-information-entropy], compression [kolmogorov1963, rissanen1978] and even fractal dimension [barnsley1993a]. Physicist Leonard Susskind believes complexity may be the key to a unified theory of physics [gefter2014, susskind2014]. Cyberneticist Francis Heylighen recently argued that goals are attractors of dynamical systems that self organise in complex reaction networks [heylighen2023b]. If complex reaction networks do self organise as he argues [heylighen2008], then it goes some way towards explaining the origins of goal directed behaviour, and thus life. Finally, many hold that complexity is the key to general intelligence. Language models like GPT-4 amount to compressed representations of human language [deletang2024]. Simpler objects can be compressed to greater extents, because they exhibit self similarity. As a result, some hold that compression is general intelligence, meaning a general reinforcement learning agent like AIXI [hutter2010] can use Solomonoff Induction [solomonoff1978] to maximise expected reward across a wide range of environments. Yet for all that complexity seems to be at the heart of every matter, it has profound flaws. As a qualitative indicator of how subjectively difficult a system might be to understand, it makes perfect sense. It makes far less as an indicator of anything objective. Ockham’s Razor is the epistemic principle that simpler statements are more likely to hold true. It can be understood as the claim that our subjective perceptions of complexity reflect an objective property of our environment. There is no obvious reason this should be the case, and yet it is [sober2015]. Simpler statements tend to be more accurate representations of reality. The aforementioned Solomonoff Induction formalises Ockham’s Razor, meaning AIXI is based on the premise that simpler models are more accurate representations of the environment. AIXI is a superintelligence in the sense that it maximises Legg-Hutter intelligence [legg2007]. However, Jan Leike later showed that any claim regarding AIXI’s performance is “entirely subjective”. Legg-Hutter intelligence is measured with respect to a fixed Universal Turing Machine (UTM), and AIXI is only optimal if it uses exactly the same UTM [leike2015]. This calls into question the viability of complexity based induction systems in interactive settings. Their performance is subjective, and from a pragmatic standpoint it is only objective performance claims that matter. Leike’s result suggests there could be no correlation between objective performance and subjective complexity. This concurs with what seems intuitively obvious, that complexity is an aspect of interpretation. Complexity is a measure of form, not function. So why does the subjective perception of simplicity tend to correlate with objective performance?
@@ 1.0.1 What exactly is complexity supposed to indicate?:
As it is used in AGI research [hutter2010], complexity is intended to help us infer the program which generated or caused data. If one can identify that which caused past data, then one can “generalise” to predict the outcomes in future interactions, to maximise performance [bennett2023c]. We are concerned with adaptation or “the ability to generalise” [chollet2019], not any specific circumstance. This is because any system can eventually identify cause given enough data and memory (by simply rote learning every outcome until it has a complete behavioural specification of the causal program). So assuming one can correctly infer cause, then the amount of data one requires to do so is the sole measure of performance. We refer to this as sample efficiency. The more sample efficiently one can infer cause, the greater one’s ability to generalise and adapt to any desired end. Thus, we take intelligence to be a measure of the sample efficiency in generalisation.
@@ 1.0.2 Key questions:
We build upon previous work [bennett2023b, bennett2023c], in which:
@@
    Maximising simplicity of policies was proven unnecessary and insufficient to maximise sample efficiency [bennett2023b, prop. 3].
@@
    Maximising policy “weakness” was proven necessary and sufficient to maximise sample efficiency [bennett2023b, prop. 1,2] and identify cause [bennett2023c]. In experiments, weak policies outperformed simple by $110-500\%$.
@@ 2 The Formalism
The following definitions are shared with [bennett2024c, bennett2023a], which apply them to biological and philosophical perspectives.
@@
  • We assume a set $\Phi$ whose elements we call states.
@@
  • We say a statement is true given a state iff it is an aspect realised by that state.
@@ 2.0.1 Learning and inference definitions:
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
@@
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
@@ 3 Arguments and Results
The distinction between software and hardware is unsuitable for reasoning about cause [bennett2023c]. The performance of a software agent in an interactive setting is subjective [leike2015], as its behaviour depends on hardware which interprets it. Hardware is an “abstraction layer” between software and the surrounding environment. If we are to understand complexity, we must understand what the concept entails in the absence of such abstraction layers. We must ascertain what is objective rather than subjective, and so we must begin at the level of the environment rather than the agent. To this end, previous work [bennett2023b, bennett2023c, bennett2023d] proposed a “de facto” pancomputational [piccinini2015] model of all conceivable environments and aspects thereof, which we refine and extend. While the formalism used here is a departure from past work, it is equivalent with respect to those previously published proofs we reference [bennett2023b]. The formalism is not computational in the sense of relying on symbols, quantities, or any other high level abstraction interpreted by a human mind. The assumptions we make are extremely weak.
Axiom 1: When there are things, we call these things the environment.
Axiom 2: The environment has at least one state [^4] Or, if the reader prefers, there are as many environments as there are states.. If there is more than one state, then there is at least one “dimension” along which things can differ.
A dimension is a set of points, for example time. We do not make the additional claim that such sets of points must be ordered. Each state is the environment at a different point in one or more dimensions. States are “extended” along dimensions, meaning no two states can occupy the same points in all dimensions (in other words, states are like the environment perceived from different positions in time or some other dimensions by an omniscient observer). These are all the assumptions we need for our version of pancomputationalism. We don’t even need to speculate about any internal structure states might have, or what dimensions might be. Any “fact” about a state’s internal structure can be defined by its relation to other states. The existence of sets is implied by the existence of states (because there is more than one “thing” that exists), and a fact is just the set of states in which it holds (the truth conditions of a “declarative program”).
@@
As truth is defined in existential terms, there is nothing which is not a fact of the environment, and no environment which we cannot describe as a set of facts. In other words, this is a minimalist formalism of everything. An environment could be like our own, or not. It could be deterministic, in which case states follow a sequence. It could be non-deterministic, in which case they don’t [^5] In any case, the difference between deterministic and non-deterministic seems meaningless when you consider that DFAs and NFAs are equivalent [rabinscott1959].. It could be fantastical with magic and true names. It could even be a world constructed through the subjective experience of its inhabitants. All that matters is that an environment has states, and from the relations between them we obtain the set of all facts. Facts as relations between states let us avoid anything like a universal set, because states are otherwise irreducible. An aspect of a state is just a set of facts about that state. With aspects in hand, we can define abstraction. An aspect is akin to a logical statement. It has a truth value given a state. We also define its extension (def. 2), which is all other aspects of which it is a part. This implies a heirarchy or “lattice” of aspects. Intuitively, an abstraction layer is like a window through which one can view part of the environment. A laptop computer could function as an abstraction layer, as could all or part of the system in which an embodied and embedded organism enacts cognition [thompson2007]. In precise terms an abstraction layer is implied by a vocabulary, which is a set of declarative programs. A vocabulary implies a formal language whose rules are determined by relations between states. An abstraction layer implies a set of “$\mathfrak{v}$-tasks” (def. 3), each of which is behaviour that defines a system. The formalisation of policies as causal intermediaries between inputs and outputs [putnam1967] then develops this into a causal depiction of goal directed behaviour (assuming a policy is implied by the inputs and outputs). For example, an organism could be a policy for a $\mathfrak{v}$-task which is behaviour that organism might enact. Again, we must emphasise this is a first principles approach. We do not assume symbols and Turing machines. Inputs, outputs and policies are all just sets of declarative programs. Whether something is goal directed is determined by the relations between states. The environment makes only one sort of value judgement (existence or non-existence), and is otherwise impartial. Goal directed behaviour is a value judgement, so we formalise this impartiality as a uniform distribution over tasks. Finally, we must now define complexity. The claims we make in the rest of this paper pertain to this notion of complexity.
@@
Complexity of Extension: If two statements have the same extension, then the simpler is the smaller set (def. 5). The complexity of an extension is the simplest statement of which it is the extension. This is like other formal notions of complexity [kolmogorov1963, rissanen1978], but facilitates comparison of abstraction layers.
@@ 3.1 Implications for Complexity
In accord with definition 2, the absence of abstraction means the vocabulary is the set of all declarative programs, meaning $\mathfrak{v}=P$. It follows that for every $l\in L_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ such that $\bigcap l=f$. Statements $l$ and $\{f\}$ are equivalent iff $E_{l}=E_{\{f\}}$, which is exactly the case here because $\bigcap l=f$. [bennett2023b, prop. 1,2] shows maximising weakness is necessary and sufficient to maximise the probability of generalisation, which means weakness maximises sample efficiency (is the optimal proxy). This means sample efficiency is determined by the cardinality of extension. For every correct policy $l$ of every task in $\Gamma_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$. Policy complexity can be minimised regardless weakness, because the simplest representation of every extension is simplicity $1$.∎
In this sense, complexity is an illusion created by abstraction. In the absence of any particular abstraction, all behaviours (extensions) are implied by statements of the same complexity. To be clear, we are not repeating the claim made by others [leike2015] that if the interpreter used by a complexity based induction system matches one used to to compute an objective value for complexity, then that induction system will be optimal in the sense of eventually learning the correct policy [^6] To quote verbatim: “Legg-Hutter intelligence is measured with respect to a fixed UTM. AIXI is the most intelligent policy if it uses the same UTM.” [leike2015]. We are claiming that if interpretation is truly objective, then $\mathfrak{v}=P$ and complexity has absolutely nothing to do with intelligence [^7] Intelligence here meaning not just eventual generalisation, but the efficiency thereof.. There is no objective notion of complexity. However, when we take empirical measurements it is inevitably through an abstraction layer, for which $\mathfrak{v}\neq P$. In that context simpler forms have been observed to generalise more efficiently. This raises the question; what additional assumptions can we make that would explain the correlation?
@@ 3.1.1 Time, space and causal confounding:
We now make the additional assumption that vocabularies are finite. Every aspect of the world in which we exist appears to be spatially extended, meaning no two things occupy the same space at the same time. For the sake of understanding complexity we assume this is true of all environments. We hold that this justifies the assumption of a finite vocabulary, because in our spatially extended environment the amount of information which can influence a system is finite, within a finite area [bekenstein1981].
@@
We already have that policy weakness causes sample efficiency, in that it is necessary and sufficient to maximise it in order to maximise sample efficiency. Continuing from proof 1, in a finite vocabulary, there may not exist $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$, which means the complexity of all extensions will not be the same. If we choose any vocabulary in which weaker aspects take simpler forms, then simplicity will be correlated with weakness and so will also be correlated with sample efficiency. This means we would choose $\mathfrak{v}$ s.t. for all $a,b\in L_{\mathfrak{v}}$, the simpler statement has the larger extension, meaning $a<_{w}b\leftrightarrow a<_{d}b$. For example, suppose $P=\{a,b,c…\},\ a=\{1,2,4\},\ b=\{1,3,4\},\ \mathfrak{v}=\{a,b\},\ L_{\mathfrak{v}}=\{\{a\},\{b\},\{a,b\}\}$, then it follows $\{a,b\}<_{w}\{a\},\ \{a,b\}<_{w}\{b\},\ \{a,b\}<_{d}\{a\},\ \{a,b\}<_{d}\{b\}$. ∎